# VideoChat-Flash: Hierarchical Compression for Long-Context Video Modeling

### v1

@@ Abstract
Long-context modeling is a critical capability for multimodal large language models (MLLMs), enabling them to process long-form contents with implicit memorization. Despite its advances, handling extremely long videos remains challenging due to the difficulty in maintaining crucial features over extended sequences. This paper introduces a Hierarchical visual token Compression (HiCo) method designed for high-fidelity representation and a practical context modeling system VideoChat-Flash tailored for multimodal long-sequence processing. HiCo capitalizes on the redundancy of visual information in long videos to compress long video context from the clip-level to the video-level, reducing the compute significantly while preserving essential details. VideoChat-Flash features a multi-stage short-to-long learning scheme, a rich dataset of real-world long videos named LongVid, and an upgraded “Needle-In-A-video-Haystack” (NIAH) for evaluating context capacities. In extensive experiments, VideoChat-Flash shows the leading performance on both mainstream long and short video benchmarks at the 7B model scale. It firstly gets 99.1% accuracy over 10,000 frames in NIAH among open-source models.
@@ 3.1 HiCo: Efficient Long Video Modeling
For video MLLMs, the ability to perceive timestamps is also a crucial capability. Unlike previous works [40, 59] that rely on additional modules or designs to achieve this, we employ a simple timestamp-aware text prompt before the video context: “The video lasts for N seconds, and T frames are uniformly sampled from it.” We find that this straightforward approach is sufficient to enable the model to perceive the timestamps of the input video, achieving excellent performance on temporally sensitive tasks such as temporal grounding. (see Tab. 3 for details)
@@ 4 Experiments
Model FLOPs Memory(G) VideoMME
(T) Train Infer Long
@@
LongVILA [57] 455.1 56.7 18.1 47.0
LongVA [65] 316.8 50.6 17.4 46.2
VideoChat-Flash (ours) 29.6 35.2 15.5 53.9
@@
LongVILA [57] 2946.3 oom 26.8 47.0
LongVA [65] 1870.1 oom 23.7 46.2
VideoChat-Flash (ours) 113.3 40.2 16.1 53.9
@@
LongVILA [57] 28707.7 oom 79.7 47.0
LongVA [65] 16583.7 oom 48.3 46.2
VideoChat-Flash (ours) 519.7 59.4 18.5 53.9
@@
LongVILA [57] 2368810.0 oom oom 47.0
LongVA [65] 1289490.0 oom oom 46.2
VideoChat-Flash (ours) 15786.0 oom 47.4 53.9
Table 1: Comparison of FLOPs and Cuda memory. FLOPs and memory are estimated by the context length of the input LLM
@@ 4.1 General Video Understanding Evaluation
Model Size #tokens MVBench PerceptionTest LongVideoBench MLVU VideoMME (w/o & w sub.) LVBench per frame Avg Val Val M-Avg Overall Long Avg Avg. Duration 16s 23s 473s 651s 1010s 2386s 4101s Proprietary Models GPT4-V [35] - - 43.7 - 59.1 49.2 59.9/63.3 53.5/56.9 - GPT4-o [36] - - 64.6 - 66.7 64.6 71.9/77.2 65.3/72.1 30.8 Gemini-1.5-Pro [39] - - 60.5 - 64.0 - 75.0/81.3 67.4/77.4 33.1 Open-Source MLLMs VideoLLaMA2 [10] 7B 72 54.6 51.4 - 48.5 47.9/50.3 42.1/43.8 - VideoLLaMA2 [10] 72B 72 62.0 57.5 - - 62.4/64.7 57.6/59.0 - VideoChat2-HD [26] 7B 72 62.3 - - 47.9 45.3/55.7 39.8/53.9 - InternVideo2-HD [50] 7B 72 67.2 63.4 - - 49.4/ - - - IXComposer-2.5 [64] 7B 400 69.1 34.4 - 37.3 55.8/58.8 - - InternVL2 [9] 8B 256 65.8 - 54.6 64.0 54.0/56.9 - - InternVL2 [9] 76B 256 69.6 - 61.1 69.9 61.2/62.8 - - InternVL2.5 [8] 8B 256 72.0 - 60.0 68.9 64.2/66.9 - - Qwen2-VL [46] 7B - 67.0 62.3 - - 63.3/69.0 - - Qwen2-VL [46] 72B - 73.6 68.0 - - 71.2/77.8 - 41.8 LLaVA-NeXT-Video [67] 7B 144 53.1 48.8 49.1 - - /46.5 - - LLaVA-OneVision [23] 7B 196 56.7 57.1 56.3 64.7 58.2/61.5 - - LLaVA-OneVision [23] 72B 196 59.4 66.9 61.3 68.0 66.2/69.5 - - LLaVA-Video [68] 7B 676 58.6 67.9 58.2 70.8 63.3/69.7 - - Open-Source Long Video MLLMs LLaMA-VID [28] 7B 2 41.9 44.6 - 33.2 25.9/ - - 23.9 Kangaroo [32] 8B 256 61.0 - 54.8 61.0 56.0 / 57.6 46.7 / 49.3 39.4 LongVILA [57] 7B 196 - - - - 57.5/60.6 47.0/52.1 - LongVA [65] 7B 144 - - - 56.3 52.6/54.3 46.2/47.6 - LongLLaVA [48] 9B 144 49.1 - - - 43.7/ - 36.4/ - - LongVU [41] 7B 64 66.9 - - 65.4 - /60.6 - /59.5 - VideoChat-Flash @224 7B 16 73.2 75.6 64.2 74.5 63.0/69.0 52.8/62.1 47.2 VideoChat-Flash @448 7B 16 73.9 76.2 64.7 74.5 64.7/69.8 53.9/63.2 48.2
@@
As shown in Tab. 2, our VideoChat-Flash gives the best results on MVBench (73.7%) and Perception Test (76.2%) among all, surpassing other approaches by a notable margin. We have achieved new state-of-the-art (SOTA) results among 7B models. Moreover, our performance even surpasses that of models with much larger scales, such as InternVL2-76B and VideoLLaMA2-72B, and proprietary models like GPT4o and Gemini-1.5-Pro. This demonstrates the effectiveness of our hierarchical compression (HiCo) in preserving crucial temporal variations as well as our dynamic sampling strategy.
@@
Tab. 2 gives that VideoChat-Flash not only outperforms previous open-source general MLLMs of the same model scale but also surpasses a host of MLLMs designed specifically for long videos on the long video understanding benchmarks. Especially, it gets the new SOTA performance (74.5%) on MLVU, even compared with GPT4-V and GPT4-o. This empirically proves that VideoChat-Flash with HiCo can handle long multimodal contexts effectively. More importantly, despite performance, the training and inference efficiency of our model is also much higher than that of previous ones. As in Tab. 1, even when processing short videos, the compute load of our model is only one-tenth that of previous models. Meanwhile, as the number of input frames increases, the difference becomes more and more pronounced. On a single A100 device, only our model can complete the inference on 10,000 frames (approximately equivalent to a 3-hour video). Concretely, VideoChat-Flash’s compute load is two orders of magnitude lower than that of LongVILA [57] (15,786 vs. 2,368,810 TFLOPs).
@@
Table 3: Results on temporal grounding task.

### v2

@@ Abstract
Long-context modeling is a critical capability for multimodal large language models (MLLMs), enabling them to process long-form contents with implicit memorization. Despite its advances, handling extremely long videos remains challenging due to the difficulty in maintaining crucial features over extended sequences. This paper introduces a Hierarchical visual token Compression (HiCo) method designed for high-fidelity representation and a practical context modeling system VideoChat-Flash tailored for multimodal long-sequence processing. HiCo capitalizes on the redundancy of visual information in long videos to compress long video context from the clip-level to the video-level, reducing the compute significantly while preserving essential details. VideoChat-Flash features a multi-stage short-to-long learning scheme, a rich dataset of real-world long videos named LongVid, and an upgraded “Needle-In-A-video-Haystack” (NIAH) for evaluating context capacities. In extensive experiments, VideoChat-Flash shows the leading performance on both mainstream long and short video benchmarks at the 2B and 7B model scale. It firstly gets 99.1% accuracy over 10,000 frames in NIAH among open-source models.
@@ 3.1 HiCo: Efficient Long Video Modeling
For video MLLMs, the ability to perceive timestamps is also a crucial capability. Unlike previous works [48, 69] that rely on additional modules or designs to achieve this, we employ a simple timestamp-aware text prompt after the video context: “The video lasts for N seconds, and T frames are uniformly sampled from it.” We find that this straightforward approach is sufficient to enable the model to perceive the timestamps of the input video, achieving excellent performance on temporally sensitive tasks such as temporal grounding. (see Tab. 3 for details)
@@ 4 Experiments
Model FLOPs Memory(G)
(T) Train Infer
@@
LongVILA [66] 455.1 56.7 18.1
LongVA [75] 316.8 50.6 17.4
VideoChat-Flash (ours) 29.6 35.2 15.5
@@
LongVILA [66] 2946.3 oom 26.8
LongVA [75] 1870.1 oom 23.7
VideoChat-Flash (ours) 113.3 40.2 16.1
@@
LongVILA [66] 28707.7 oom 79.7
LongVA [75] 16583.7 oom 48.3
VideoChat-Flash (ours) 519.7 59.4 18.5
@@
LongVILA [66] 2368810.0 oom oom
LongVA [75] 1289490.0 oom oom
VideoChat-Flash (ours) 15786.0 oom 47.4
Table 1: Comparison of FLOPs and Cuda memory. FLOPs and memory are estimated by the context length of the input LLM.
@@ 4.1 General Video Understanding Evaluation
Model Size #tokens MVBench PerceptionTest LongVideoBench MLVU VideoMME (w/o & w sub.) LVBench per frame Avg Val Val M-Avg Overall Long Avg Avg. Duration 16s 23s 473s 651s 1010s 2386s 4101s Proprietary Models GPT4-V [43] - - 43.7 - 59.1 49.2 59.9/63.3 53.5/56.9 - GPT4-o [44] - - 64.6 - 66.7 64.6 71.9/77.2 65.3/72.1 30.8 Gemini-1.5-Pro [47] - - 60.5 - 64.0 - 75.0/81.3 67.4/77.4 33.1 Small Size MLLMs Qwen2-VL [55] 2B - 63.2 - - - 55.6/60.4 - - InternVL2.5 [11] 2B 256 68.8 - 46.0 61.4 51.9/54.1 - - VideoChat-Flash @448 2B 16 70.0 70.5 58.3 65.7 57.0/63.9 44.9/54.0 42.9 Open-Source MLLMs VideoLLaMA2 [13] 7B 72 54.6 51.4 - 48.5 47.9/50.3 42.1/43.8 - VideoLLaMA2 [13] 72B 72 62.0 57.5 - - 62.4/64.7 57.6/59.0 - VideoChat2-HD [31] 7B 72 62.3 - - 47.9 45.3/55.7 39.8/53.9 - InternVideo2-HD [59] 7B 72 67.2 63.4 - - 49.4/ - - - IXComposer-2.5 [74] 7B 400 69.1 34.4 - 37.3 55.8/58.8 - - InternVL2 [12] 8B 256 65.8 - 54.6 64.0 54.0/56.9 - - InternVL2 [12] 76B 256 69.6 - 61.1 69.9 61.2/62.8 - - InternVL2.5 [11] 8B 256 72.0 - 60.0 68.9 64.2/66.9 - - Qwen2-VL [55] 7B - 67.0 62.3 - - 63.3/69.0 - - Qwen2-VL [55] 72B - 73.6 68.0 - - 71.2/77.8 - 41.8 LLaVA-NeXT-Video [78] 7B 144 53.1 48.8 49.1 - - /46.5 - - LLaVA-OneVision [27] 7B 196 56.7 57.1 56.3 64.7 58.2/61.5 - - LLaVA-OneVision [27] 72B 196 59.4 66.9 61.3 68.0 66.2/69.5 - - LLaVA-Video [79] 7B 676 58.6 67.9 58.2 70.8 63.3/69.7 - - Open-Source Long Video MLLMs LLaMA-VID [33] 7B 2 41.9 44.6 - 33.2 25.9/ - - 23.9 Kangaroo [38] 8B 256 61.0 - 54.8 61.0 56.0 / 57.6 46.7 / 49.3 39.4 LongVILA [66] 7B 196 67.1 58.1 57.1 - 60.1/65.6 47.0/52.1 - LongVA [75] 7B 144 - - - 56.3 52.6/54.3 46.2/47.6 - LongLLaVA [57] 9B 144 49.1 - - - 43.7/ - 36.4/ - - LongVU [50] 7B 64 66.9 - - 65.4 - /60.6 - /59.5 - VideoChat-Flash @224 7B 16 73.2 75.6 64.2 74.5 64.0/69.4 53.6/61.9 47.2 VideoChat-Flash @448 7B 16 74.0 76.2 64.7 74.6 65.3/69.7 55.4/63.3 48.2
@@
As shown in Tab. 2, our VideoChat-Flash gives the best results on MVBench (74.0%) and Perception Test (76.2%) among all, surpassing other approaches by a notable margin. We have achieved new state-of-the-art (SOTA) results among 7B models. Moreover, our performance even surpasses that of models with much larger scales, such as InternVL2-76B and VideoLLaMA2-72B, and proprietary models like GPT4o and Gemini-1.5-Pro. Furthermore, our model achieved a comprehensive lead in results at the 2B parameter scale, even surpassing some models with 7B or even 72B parameters. This demonstrates the effectiveness of our approach in enhancing video perception capabilities.
@@
Tab. 2 gives that VideoChat-Flash not only outperforms previous open-source general MLLMs of the same model scale but also surpasses a host of MLLMs designed specifically for long videos on the long video understanding benchmarks. Especially, it gets the new SOTA performance (74.7%) on MLVU, even compared with GPT4-V and GPT4-o. This empirically proves that VideoChat-Flash with HiCo can handle long multimodal contexts effectively. More importantly, despite performance, the training and inference efficiency of our model is also much higher than that of previous ones. As in Tab. 1, even when processing short videos, the compute load of our model is only one-tenth that of previous models. Meanwhile, as the number of input frames increases, the difference becomes more and more pronounced. On a single A100 device, only our model can complete the inference on 10,000 frames (approximately equivalent to a 3-hour video). Concretely, VideoChat-Flash’s compute load is two orders of magnitude lower than that of LongVILA [66] (15,786 vs. 2,368,810 TFLOPs).
@@
Table 3: Results on temporal grounding task. Similar to previous Video MLLM approaches [48, 69], our model is trained without any grounding data from the Charades-STA dataset, evaluating its zero-shot capabilities.
@@ 5 Conclusions
Supplementary Material
6 Implementation Details and More Analysiss
6.1 Clip Compressor
As shown in Fig. 4, we consider four popular token compression strategies to compress the features from video clips:
  • Spatial Downsampling. Applying spatial operations (pooling [65], interpolation [78], and convolution (pixel shuffle) [12]) to each video frame for downsampling has been demonstrated in previous work [65, 39] as an effective method to reduce the number of video tokens. However, due to the lack of temporal interaction, this approach fails to leverage the relation between frames. We use pixel shuffle in our experiments.
  • Uneven Downsampling. Considering the similarities between adjacent frames, it is unnecessary to retain full details for every frame. We can apply down-sampling operations with different sizes across frames within a clip. Specifically, a lower down-sampling size is applied to the first frame, while higher down-sampling sizes are used for the remaining frames. Similar approaches have been validated in a recent study [62].
  • Spatio-Temporal Resampler. Using a learnable compressor, such as a Q-Former [31] or a cross-attention layer, to compress spatiotemporal tokens. However, this approach requires a large amount of data for effective learning. In training, we observe that the Q-Former barely converges well in our setting. So in our ablations, we adopt a single-layer cross-attention instead.
  • Similar Token Merging. We directly merge similar tokens, using the ToMe [5] approach.
6.2 Visual Dropout in LLM
Figure 8: Visual redundancy in long context across layers. We conduct experiments on Qwen2-7B (28 layers) and test the impact of droping 50% of the visual tokens from shallow to deep layers.
Visual token redundancy in LLM inference.
As shown in Fig. 8, we find that even when half of the tokens are discarded at the shallow layers of the LLM, the performance of long video understanding only degrades marginally. This indicates that despite high compression at the clip level (encoding each frame into only 16 tokens), there remains considerable redundancies between clips when their representations are interacted in the LLM. Furthermore, we find the overall understanding performance gets better as the dropout happens in the deeper layer of the model. Remarkably, at approximately two-thirds of the LLM’s depth, the performance even surpasses that of the no-discard baseline. This might suggest that in the deeper layers of the network, an excess of visual tokens may interfere with the model’s reasoning process. For the drop type, we observe that uniform drop often outperforms attention-based selection in the shallow layers. We suppose, at these layers, the LLM has not yet fully determined the specific locations to focus on. As a result, relying on attention may introduce bias.
Visualization of visual attention map.
As shown in the  Fig. 9, for long video context, the attention of text tokens is relatively dispersed in the shallow layers of the network. However, as the layers deepen, the attention gradually becomes focused on specific regions. Thus, we believe that the attention scores in the deeper layers are more reliable, while those in the shallow layers may be prone to bias.
Figure 9: Visualization of the attention scores from the last textual token to visual tokens at each layer of the network.
7 Training Details
7.1 Training hyperparameters.
As shown in Table 1, the training details and hyperparameters for each stage of our VideoChat-Flash model are presented.
Stage-1 Stage-2 Stage-3 Stage-4
Vision Resolution×$\times$×Num. frames 224 224 $\times$8 224$\times$(64$\sim$512) 224$\times$(64$\sim$512)
#Tokens 16$\times$4 16$\times$8 16$\times$(64$\sim$512) 16$\times$(64$\sim$512)
Data Dataset Image & Short Video Image & Short Video (Multi)-Image & Short/Long Video (Multi)-Image & Short/Long Video
#Samples 1M 4M 3.2M 0.3M
Model Trainable Projector Full Model Full Model ViT&Projector
7.6B LLM 20.0M 7.9B 7.9B 0.3B
Training Batch Size 512 256 256 256
LR of vision encoder 1×$\times 10^{-3}$ 2×$\times 10^{-6}$ 2×$\times 10^{-6}$ 2×$\times 10^{-6}$
LR of connector & LLM 1×$\times 10^{-3}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$ 1 ×$\times 10^{-5}$
Epoch 1 1 1 1
Table 9: Training details of each training stage for the VideoChat-Flash-7B model
7.2 Training Data
Stage 1: Video-Language Alignment.
In this stage, we use 558k image-text pairs from LCS-558K [37] and 481k short video-text pairs from S-MiT [42].
Stage 2: Short Video Pre-training.
To enhance the model’s understanding of visual concepts, we conduct visual pre-training using 3.5 million images and 2.5 million short video-text pairs.
  • Video Description Data. We utilize the video description data recaptioned with VideoChat2 [31] from WebVid2M [3].
  • Detailed Video Description Data. We employ the 323k detailed video description data recaptioned with Gemini [47] from WebVid [3] and Kinetics [24], as in previous work [49].
  • Detailed Image Description Data. We use the 3.5 million detailed image description data recaptioned with LLava-NeXT-34B [78] from the following datasets: COCO118K, BLIP558K, and CC3M, as provided by previous work [27].
  • Text Data. To enhance the model’s language understanding capabilities, we incorporate 143K samples from the Evo-Instruct dataset [7].
Stage 3: Joint Short & Long Video Instruction tuning.
To enable the model to handle a wide variety of video tasks, we collect 3.5 million instruction fine-tuning samples, including 1.1M images, 1.7M short videos (under 60 seconds), and 0.7M long videos (60$\sim$3600 seconds).
  • Image Instruction data. We primarily utilized single-image instruction data from LLava-NeXT [78], Allava [7], and ShareGPT4-o [12, 59]. Additionally, we incorporated multi-image data provided by LLaVA-Interleave [28].
  • Short Video Instruction data. We primarily utilized short video data from VideoChat2 [31] and InternVideo2 [59] for instruction fine-tuning. Additionally, we incorporated data annotated with GPT4-o from previous works, including ShareGPT4o [12, 59], VideoChatGPT-Plus [39], LLaVA-Video-178K [79] and LLava-Hound [76].
  • Long Video Instruction data. We primarily utilized long video instruction data from MoiveChat [52], Vript [68] and our LongVid.
8 Dataset Details of LongVid
The videos of LongVid are curated from 4 open-source video datasets: Ego4D [18], HowTo100M [41], HD-VILA [67], and MiraData [23]. We provide details of the data construction pipeline for each dataset as follows.
8.1 Ego4D
For ego-centric videos, we adopt 3,662 long videos from the Ego4d [18] and leverage Ego4DHcap [22] as the corresponding captions. Ego4DHcap gives hierarchical captions for short, medium, and long video segments. For the short video captioning task, we directly utilize these captions, while for the dense caption task, we concatenate captions in the lower level to form a dense one. For example, we merge all short video captions in a medium video segment to create a dense medium-level one, and the dense caption of long video segments can be formed by concatenating multiple medium-level video captions.
We also build event relation recognition and temporal grounding tasks based on captions of short video segments. For the event relation recognition task, models are required to choose the right order of an event sequence. Since we find the captions of short videos are highly concise and event-oriented, we use them as the event labels and serially put the short captions in a medium-level video segment as the ground-truth event relationship. For the temporal grounding task, we use the short video captions with the corresponding timestamps as the ground-truth, and randomly select other timestamps in the current medium video segments as the false options.
8.2 MiraData
MiraData [23] provides multi-level captions for large-scale minute-level movie segments. Apart from short and dense captions that are used for short and dense video captioning tasks, it also provides multiple fine-grained captions that focus on various specific perspectives, such as the main object, background, camera movements, and video style. We use an open-source LLM (Qwen-72b [2]) to extract the event and background labels from the main object and background captions, respectively, and we put the labels of a long video in the right order as the ground truth of the event/background relation recognition task. For the temporal grounding task, we use the event label with the corresponding timestamp as the ground-truth option.
8.3 HowTo100M
HowTo100M [41] includes more than 1 million long-duration how-to videos. We adopt HowToInterlink7M [54], a video captioning dataset that provides refined interleaved video captions of HowTo100M videos as short and dense video captions. For the event relationship recognition and temporal grounding tasks, we use HTStep [1], a large-scale dataset containing temporal annotations of instructional steps in HowTo100M videos.
8.4 HD-VILA
While previous datasets focus on long videos in specific domains, we also select part of the videos from HD-VILA [67], a large-scale video dataset that includes various in-the-wild videos. We argue that adding these videos into training could enhance the model’s ability to process long videos in some uncommon domains. For HD-VILA videos, we adopt the captions of Panda-70M [10]. Specifically, we filter consecutive video segments that can be re-constructed into more than 60s long videos from the 10M training subset and utilize these captions as the video short/dense captioning and temporal grounding tasks. The event labels are also extracted from these captions in the same way as MiraData [23].